# On the optimal prediction of extreme events in heavy-tailed time series with applications to solar flare forecasting

### v1

@@ 5 Data Analysis
Table 5.1: A summary of the performance of various predictors. For each pair of $h$ and $p$, the goal was to predict exceedance of the $p$th marginal quantile $h$ steps ahead. In each row, for each metric, the largest value is in boldface. For more details, see Table A.2.
@@ 6 Acknowledgements
We thank Dr. Janet Machol and Prof. Ward Manchester for advice on processing the GOES data. We acknowledge NOAA and the NCEI for making that data available.
@@ A.1.3 ROC Dominance Property
(i) Let $\tau_{1},\tau_{2}\in\mathbb{R}$ such that $\tau_{1}\leq\tau_{2}$. Since tail functions are monotone non-increasing,
@@
Thus, $\text{TPR}^{*}(\cdot)$ and $\text{FPR}^{*}(\cdot)$ are monotone non-increasing. In a similar way, $\text{TPR}(\cdot)$ and $\text{FPR}(\cdot)$ can be shown to be monotone non-increasing.
@@ A.2.3 A Consistency Lemma for Autoregressive Models
A.2.4 Regular Variation
@@
Apply Corollary 2.1.14 of Kulik & Soulier (2020). ∎
A.2.5 Proof of Lemma 3.3
@@
A.2.6 Proof of Lemma 3.4
Proof.
Let
Set $\delta=\delta_{a}\vee\delta_{b}$. Since $\delta/2\leq 1$,
It follows that
where $\nu_{\epsilon}$ is defined by
$\nu_{\epsilon}(dt)=\kappa_{+}(t)\alpha|t|^{-\alpha-1}\,dt.$
$\lambda(X,Y)=\lim_{n\to\infty}n\mathbb{P}(X>c_{n}x_{0},Y>c_{n}y_{0})=\nu((x_{0},\infty)\times(y_{0},\infty))$
where $x_{0}$ and $y_{0}$ are real numbers such that $c_{n}x_{0}\sim F_{X}^{\leftarrow}(1-1/n)$ and $c_{n}y_{0}\sim F_{Y}^{\leftarrow}(1-1/n)$. Also, since $\nu=\sum_{j=0}^{\infty}\nu_{A_{j}\epsilon_{j}}$ it remains to compute $\nu_{A_{j}\epsilon_{j}}((x_{0},\infty)\times(y_{0},\infty))$, which will involve examining the geometry of the support of $\nu$ and considering cases.
For any $x_{0}>0$,
$=\nu_{Z}((x_{0},\infty)\times\mathbb{R})$
$=x_{0}^{-\alpha}\nu_{Z}((1,\infty)\times\mathbb{R})$
We have
[…]
If $a_{j}<0$, then we need $t<1/a_{j}$, so
[…]
Hence, whether $a_{j}$ is zero, positive, or negative,
Using (A.36), in order for $\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n})=1$, we must have
$x_{0}^{-\alpha}\sum_{j=0}^{\infty}\nu_{A_{j}\epsilon_{j}}((1,\infty)\times\mathbb{R})=1,\ \text{or}\ x_{0}^{\alpha}=\sum_{j=0}^{\infty}\kappa_{+}(a_{j})|a_{j}|^{\alpha}=\eta_{+}({a},{0}).$
Since $X$ is regularly varying and $p_{X}>0$, $\overline{F}_{X}(F_{X}^{\leftarrow}(p))\sim 1-p$ as $p\uparrow 1$ by Lemma A.4. This implies that $\lim_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n))=1$. It follows from Lemma A.5 that as $n\to\infty$, $F_{X}^{\leftarrow}(1-1/n)\sim x_{0}c_{n}$. In a similar way, we can show that when
$y_{0}^{\alpha}=\sum_{j=0}^{\infty}\kappa_{+}(b_{j})|b_{j}|^{\alpha}=\eta_{+}({b},{0}),$
then $F_{Y}^{\leftarrow}(1-1/n)\sim y_{0}c_{n}$ as $n\to\infty$.
Now,
$\lambda(X,Y)$ $=\lim_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))$
$=\lim_{n\to\infty}n\mathbb{P}(X>x_{0}c_{n},Y>y_{0}c_{n})$
$=\lim_{n\to\infty}n\mathbb{P}(c_{n}^{-1}Z\in(x_{0},\infty)\times(y_{0},\infty))$
$=\nu_{Z}((x_{0},\infty)\times(y_{0},\infty))$
$=\sum_{j=0}^{\infty}\nu_{A_{j}\epsilon_{j}}\left([(x_{0},\infty)\times(y_{0},\infty)]\cap{\text{span}}\{A_{j}\}\right).$
[…]
Similarly, if $a_{j},b_{j}<0$, then
$\nu_{A_{j}\epsilon_{j}}((x_{0},\infty)\times(y_{0},\infty))=\kappa_{+,+}(a_{j},b_{j})\left(\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|b_{j}|^{\alpha}}{\eta_{+}({b},{0})}\right),$
and the tail dependence coefficient has the specified form. ∎
@@
Let $X$ be a regularly varying random variable with tail index $\alpha$ and extremal skewness $p_{X}>0$. Let $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$ be real sequences such that as $n\to\infty$,
$a_{n}\to\infty,\ b_{n}\to\infty,\ \mathbb{P}(X>a_{n})\sim\frac{1}{n},\ \text{and}\ \mathbb{P}(X>b_{n})\sim\frac{1}{n}.$
Then $a_{n}\sim b_{n}$ as $n\to\infty$.
@@
Because $X$ is regularly varying with positive extremal skewness, there exists a slowly varying function $\ell$ such that for large enough $x$, $\overline{F}_{X}(x)=\ell(x)x^{-\alpha}$. Also, since $\overline{F}(a_{n})/\overline{F}(b_{n})\to 1$ as $n\to\infty$, $\overline{F}(a_{n_{k}})/\overline{F}(b_{n_{k}})\to 1$ as $k\to\infty$. Combining these observations with (A.38), we get
11 11 $=\lim_{k\to\infty}\frac{\overline{F}(a_{n_{k}})}{\overline{F}(b_{n_{k}})}$
$\geq\lim_{k\to\infty}\frac{\overline{F}_{X}((1-\epsilon)b_{n_{k}})}{\overline{F}_{X}(b_{n_{k}})}$
$\geq(1-\epsilon)^{-\alpha},$
@@
A.2.7 Half-Spaces in $\mathbb{R}^{d}$
Definition A.9.
Definition A.10.
Definition A.11.

### v2

@@ 4.4 Prediction Results
Table 4.1: A summary of the performance of various predictors. The AR models were fit using OLS. For each pair of $h$ and $p$, the goal was to predict exceedance of the $p$th marginal quantile $h$ steps ahead. In each row, for each metric, the largest value is in boldface. For additional results, including results for AR models fit using LAD, see Table A.3.
@@
5 Supporting Information
Additional supporting information can be found online in the supporting information tab for this article.
6 Data Availability Statement
Code for the simulations and data analysis and the data for the latter can be accessed at https://github.com/victorverma/opt_extreme_event_pred. The Solar Cycle Progression product mentioned in Section 4 is located at https://www.swpc.noaa.gov/products/solar-cycle-progression.
@@ 7 Acknowledgements
We thank Dr. Janet Machol and Prof. Ward Manchester for their advice on processing the GOES data. We acknowledge NOAA and the NCEI for making that data available. SS was partially supported by the NSF grant 2319592. YC and VV were partially supported by NSF DMS 2113397, NSF PHY 2027555, NASA Federal Award No. 80NSSC23M0192 and No. 80NSSC23M0191.
@@ A.1.3 ROC Dominance Property
Remark A.5.
The performance of a binary classifier on a dataset can be summarized by a confusion matrix like that in Table A.1; the two classes are represented by + and –, and TP, FP, FN, and TN are the counts of true positives, false positives, false negatives, and true negatives, respectively. For example, TP is the number of observations belonging to the positive class that were predicted to belong to that class.
Observed
Predicted + TP FP
FN TN
Table A.1: A confusion matrix summarizing the performance of a binary classifier on a dataset; the two classes are represented by + and –, and ${\rm TP}$, ${\rm FP}$, ${\rm FN}$, and ${\rm TN}$ are the counts of true positives, false positives, false negatives, and true negatives, respectively.
The TPR, FPR, and precision are typically defined in terms of the confusion matrix as follows (Fawcett 2006):
$\text{TPR}:=\frac{\text{TP}}{TP+FN},\ \text{FPR}:=\frac{FP}{FP+TN},\ \text{precision}:=\frac{TP}{TP+FP}.$ (A.15)
In (A.15), the TPR is the sample proportion of positive observations predicted to be positive, the FPR is the sample proportion of negative observations predicted to be positive, and the precision is the sample proportion of observations predicted to be positive that actually are positive. The probabilities in Definition A.1 are the population analogues of the TPR and FPR in (A.15) when an observation $(X,Y)$ belongs to the positive class if $Y>y_{0}$ and the negative class otherwise. Similarly, the precision as defined in Definition 2.1 is the population analogue of the sample precision in (A.15).
@@
(i) Let $\tau_{1},\tau_{2}\in\mathbb{R}$ such that $\tau_{1}\leq\tau_{2}$. Since tail functions are decreasing,
@@
Thus, $\text{TPR}^{*}(\cdot)$ and $\text{FPR}^{*}(\cdot)$ are decreasing. In a similar way, $\text{TPR}(\cdot)$ and $\text{FPR}(\cdot)$ can be shown to be decreasing.
@@ A.2.3 A Consistency Lemma for Autoregressive Models
A.2.4 Proof of Proposition 3.1
Proof.
$|\widehat{U}_{t+h}-U_{t+h}^{*}|\leq|\widehat{U}_{t+h}-\widetilde{U}_{t+h}|+|\widetilde{U}_{t+h}-U_{t+h}^{*}|.$ (A.33)
$f_{j}(x_{0},x_{1},\cdots):=\sum_{k=0}^{\infty}a_{k}(\phi)x_{j+k},$
where $Y_{t}=\sum_{k=0}^{\infty}a_{k}(\phi)\epsilon_{t-k}$ is the causal moving average representation of the one-dimensional AR$(d)$ time-series. Now, the ergodicity of the map $T$ entails the ergodicity of the process $\{Z_{t}=Z_{0}\circ T^{t}\}$. In order to use Theorem 3.1 from Adams & Nobel (2012), we must also have that the collection of all half-spaces $H_{\varphi,\,\tau}$ is a separable Vapnik-Chervonenkis class; this is shown in Lemma A.3. The latter uniform law of large numbers theorem now yields
which shows that the first term on the right side in (A.33) vanishes.
Now we handle the second term on the right side in (A.33). Recall that $Z$ in (3.15) is chosen to be independent from $(Y_{t:(t-d+1)},\widehat{\phi})$. Thus, $Z$ is independent from $\widehat{Y}_{t+h}(\widehat{\phi})=\widehat{\phi}(h)^{\top}Y_{t:(t-d+1)}$ and $\widehat{\phi}(h)$ and in view of (3.15), we can write:
$\widetilde{U}_{t+h}-U_{t+h}^{*}=\mathbb{E}\left[\mathbb{I}(\widehat{\phi}(h)^{\top}Z\leq\widehat{Y}_{t+h}(\widehat{\phi}))-\mathbb{I}(\phi(h)^{\top}Z\leq\widehat{Y}_{t+h}(\phi))\Big{|}\widehat{\phi},Y_{t:(t-d+1)}\right]\$ (A.34)
Introduce
In view of (A.34),
$\leq\mathbb{E}|\mathbb{I}[\xi_{n}\leq 0]-\mathbb{I}[\xi\leq 0]|$
[…]
This completes the proof. ∎
A.2.5 Half-Spaces in $\mathbb{R}^{d}$ Form a Separable Vapnik-Chervonenkis Class
Definition A.5.
Definition A.6.
Definition A.7.
Lemma A.3.
Proof.
$\geq\|\varphi^{(k)}-\varphi\|^{-1/2},$
A.2.6 A Simulation Study for an Autoregressive Model
Optimal extreme event prediction for autoregressive models was discussed in Section 3.1; prediction when the AR coefficients are unknown was discussed in Section 3.2. Here, we use simulated data to evaluate the performance of various predictors for an AR model. We restate the model from Sections 3.1 and 3.2 for the sake of convenience:
$Y_{t}=\sum_{i=1}^{d}\phi_{i}Y_{t-i}+\epsilon_{t},\ \ \ t\in{\mathbb{Z}}.$ (A.39)
It was shown that for a stationary process satisfying (A.39), for $h\geq 1$, the optimal predictor of $\mathbb{I}(Y_{t+h}>y_{0})$ is $\mathbb{I}(\widehat{Y}_{t+h}(\phi)>\tau)$ for an appropriate choice of $\tau$ (Theorem 3.2). In Section 3.2.1, we discussed prediction of $\mathbb{I}(Y_{t+h}>F_{Y}^{\leftarrow}(p))$, $F_{Y}^{\leftarrow}(p)$ being the $p$th quantile of the marginal distribution of $Y_{t}$; we considered the use of the predictor $\mathbb{I}(\widehat{Y}_{t+h}(\widehat{\phi})\geq\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p))$, which can be estimated from training data, unlike the predictor given by Theorem 3.2, $\mathbb{I}(\widehat{Y}_{t+h}(\phi)>F_{\widehat{Y}_{t+h}(\phi)}^{\leftarrow}(p))$, which we shall call the oracle predictor.
The most straightforward way to compute the estimate $\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)$ of $F_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)$ is to use an empirical quantile, as was done in Section 3.2.1. When $\{Y_{t}\}$ is a heavy-tailed process and $p$ is close to one, an alternative is to use an extreme quantile estimator based on extreme value theory (see, e.g., Weissman 1978, Drees 2003, Chavez-Demoulin & Guillou 2018, and the references therein).
We conducted a simulation study that compared the performance of the oracle predictor, a predictor based on the empirical quantile estimator, and a predictor based on an extreme quantile estimator. Data was generated from the AR(5) model
with the $\epsilon_{t}$’s being iid $t$-distributed with one degree of freedom. The corresponding AR polynomial has roots $-3\pm 1i$, $4\pm 2i$, and $2$; since these roots lie outside the closed unit disk, it follows from Proposition 13.3.2 of Brockwell & Davis (1991) that (A.40) has a unique strictly stationary, causal solution and that the solution is invertible. Set
$\phi=(0.3,0.19,-0.035,-0.01,0.0025).$
Next, one hundred times, a time series of length $10^{4}+10^{6}$ was generated from the model in (A.40); the first $10^{4}$ observations were reserved for training, and the last $10^{6}$ were reserved for testing. Inspired by Davis et al. (1992), LAD estimation was used on the training set to compute an estimate $\widehat{\phi}$ of $\phi$ for the non-oracle predictors; the order of the model was assumed to be known. Essentially, a median regression was carried out in which the response was the observation one step ahead and the covariates were the five most recent observations. Then, for the three predictors, one-step-ahead prediction was performed for each $p$; prediction worked the same way for each $p$.
$\phi^{\top}Y_{t:(t-4)}>F_{\widehat{Y}_{t+h}(\phi)}^{\leftarrow}(p).$
$\widehat{\phi}^{\top}Y_{t:(t-4)}\geq\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p),$
where $\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)$ was computed as the $p$th sample quantile of all linear combinations of the form $\widehat{\phi}^{\top}Y_{s:(s-4)}$ that could be computed from the training set.
$\widehat{\phi}^{\top}Y_{t:(t-4)}\geq\widetilde{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p),$
where $\widetilde{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)$ was computed from the training set using the following procedure. For each of ten equally-spaced values between the third quartile of the training linear combinations and the tenth largest linear combination, a Generalized Pareto (GP) model was fit to the declustered excesses above the value. The excesses were the differences between the training observations above the value and the value, and they were declustered using the intervals method introduced in Ferro & Segers (2003). The purpose of declustering was to extract excesses that could reasonably be viewed as independent, which was important since the GP model was fit using maximum likelihood estimation. The GP distribution with location parameter $\mu$, scale parameter $\sigma>0$, and shape parameter $\xi$ has distribution function
$G(y)=\begin{cases}1-\left(1+\xi\left(\frac{y-\mu}{\sigma}\right)\right)^{-1/\xi},&\xi\neq 0\\ 1-\exp\left(-\frac{y-\mu}{\sigma}\right),&\xi=0,\end{cases}$
with the support being $[\mu,\infty)$ when $\xi\geq 0$ and $[\mu,\mu-\sigma/\xi]$ when $\xi<0$; GP distributions are frequently used to model excesses above a threshold (Embrechts et al. 1997, Coles 2001).
$\widetilde{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)=\tau+\widehat{G}^{\leftarrow}\left(\frac{p-p_{0}}{1-p_{0}}\right),$
Once all predictions were made for the test set, the precision was computed as the number of exceedances that were predicted to be exceedances divided by the number of predicted exceedances. Since there were 100 training set-test set pairs, there were 100 precisions for each pair of a predictor and a value of $p$. Finally, the asymptotic precision of the optimal predictor was computed by deriving the causal representation of the solution to (A.40) and then applying Theorem 3.4 using the first $10^{6}$ terms in each series.
The results of the simulation study are displayed in Figure A.1. Orange boxplots represent the precisions of the oracle predictor. As expected, as $p$ approaches one, the oracle precision converges to the asymptotic oracle precision, which is represented by the dashed orange line. The blue boxplots represent the precisions of the predictor based on the empirical quantile estimator; this predictor exhibits strong performance, as its precisions are mostly close to the corresponding oracle precisions. However, this is not the case when $p=0.999$; most of the precisions are smaller than the corresponding oracle precisions. This is not surprising, as the training set size was $10,000$. The black boxplots represent the precisions of the predictor based on the extreme quantile estimator. Interestingly, for each $p$, all or almost all of these precisions lie well above the corresponding oracle precisions. It turned out that the extreme quantile estimates were substantially too large, so the predictor based on the extreme quantile estimator was calibrated at a higher level than the other predictors. This is not guaranteed to yield higher precisions, but it ended up doing so in this case.
Figure A.1: Results from the autoregressive model simulation study. Data was generated from a particular AR(5) model. One hundred training set-test set pairs were generated; the model was fit to the training set, and the fitted model was used to predict exceedances of high quantiles in the test set. The quantile levels used are on the horizontal axis. On each of the 100 runs, the precision was calculated from the test set predictions. The dashed orange line represents the asymptotic precision of the optimal predictor.
A.2.7 Regular Variation
@@
Apply Corollary 2.1.14 of Kulik & Soulier (2020).
A.2.8 Proof of Lemma 3.3
@@
A.2.9 Proof of Lemma 3.4
We first prove several lemmas needed in the proof of Lemma 3.4.
@@
Let $X$ be a regularly varying random variable with tail index $\alpha$ and extremal skewness $p_{X}>0$. Let $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$ be real sequences such that as $n\to\infty$, $a_{n}\to\infty$, $b_{n}\to\infty$, and $\overline{F}_{X}(a_{n})\sim\overline{F}_{X}(b_{n})$. Then $a_{n}\sim b_{n}$ as $n\to\infty$.
@@
Since $\overline{F}_{X}(a_{n})/\overline{F}_{X}(b_{n})\to 1$ as $n\to\infty$, $\overline{F}_{X}(a_{n_{k}})/\overline{F}_{X}(b_{n_{k}})\to 1$ as $k\to\infty$. Also, because $X$ is regularly varying with positive extremal skewness, $\overline{F}_{X}$ is regularly varying with index $-\alpha$ (Kulik & Soulier 2020). Combining these observations with (A.46), we get
$1=\lim_{k\to\infty}\frac{\overline{F}_{X}(a_{n_{k}})}{\overline{F}_{X}(b_{n_{k}})}\\ \geq\lim_{k\to\infty}\frac{\overline{F}_{X}((1-\epsilon)b_{n_{k}})}{\overline{F}_{X}(b_{n_{k}})}\\ =(1-\epsilon)^{-\alpha},$
@@
The final lemma gives a sufficient condition for the existence of the tail dependence coefficient, which was defined in Definition 2.2.
@@
Let $X$ and $Y$ be random variables such that both of these conditions hold:
(i) $\mathbb{P}(X>F_{X}^{\leftarrow}(p))\sim\mathbb{P}(Y>F_{Y}^{\leftarrow}(p))\sim 1-p$ as $p\uparrow 1$.
(ii) $\lim_{n\to\infty}n\mathbb{P}(X>F_{X}^{\leftarrow}(1-1/n),Y>F_{Y}^{\leftarrow}(1-1/n))=\lambda\in[0,1]$.
Then the tail dependence coefficient $\lambda(X,Y)$ exists and equals $\lambda$:
$\lambda(X,Y)=\lim_{p\uparrow 1}\mathbb{P}(X>F_{X}^{\leftarrow}(p)\mid Y>F_{Y}^{\leftarrow}(p))=\lambda.$